# EvRWKV: A Continuous Interactive RWKV Framework for Effective Event-Guided Low-Light Image Enhancement

### v2

@@ Abstract
Capturing high-quality visual content under low-light conditions remains a challenging problem due to severe noise and underexposure, which degrade the performance of downstream applications. Traditional frame-based low-light image enhancement methods often amplify noise or fail to preserve structural details. Event cameras, offering high dynamic range and microsecond temporal resolution by asynchronously capturing brightness changes, emerge as a promising complement for low-light imaging. However, existing fusion methods fail to fully exploit this synergy, either by forcing modalities into a shared representation too early or by losing vital low-level correlations through isolated processing. To address these challenges, we propose EvRWKV, a novel framework that enables continuous cross-modal interaction through dual-domain processing. Our approach incorporates a Cross-RWKV module, leveraging the Receptance Weighted Key Value (RWKV) architecture for fine-grained temporal and cross-modal fusion, and an Event Image Spectral Fusion Enhancer (EISFE) module, which jointly performs adaptive frequency-domain noise suppression and spatial-domain deformable convolution alignment. This continuous interaction maintains feature consistency from low-level textures to high-level semantics. Extensive qualitative and quantitative evaluations on real-world low-light datasets (SDE, SDSD, RELED) demonstrate that EvRWKV achieves state-of-the-art performance, effectively enhancing image quality by suppressing noise, restoring structural details, and improving visual clarity in challenging low-light conditions.
Index Terms:
Low-light image enhancement, event cameras, cross-modal fusion, RWKV, dual-domain processing.
@@ I Introduction
Capturing high-quality visual content under low-light conditions remains a critical yet challenging task in computer vision [26]. Images and videos acquired in such environments often suffer from severe noise, motion blur, and underexposure [26], which degrade downstream applications such as autonomous driving and surveillance. Traditional frame-based low-light image enhancement (LLIE) methods [23, 47, 9, 50] have made progress, but often struggle to effectively address the inherent complexities of real-world low-light scenarios. Many methods [55, 3, 49, 33, 27]can amplify existing noise while attempting to brighten the image, leading to visually unappealing results. Furthermore, approaches relying on synthetic data for training often fail to generalize well to the diverse and unpredictable nature of real-world low-light scenes.
@@
Event cameras, using bio-inspired sensors [32], are widely used in the field of computer vision due to their high dynamic range and microsecond level time resolution [43, 59], such as image restoration [17], video super-resolution [37], optical flow [38], deblurring [18, 19, 6], etc., and have also become a promising alternative to low-light imaging scenes. By asynchronously capturing per-pixel brightness changes, events excel at preserving structural details even in extreme darkness and are less susceptible to motion blur. This makes them particularly attractive for low-light image enhancement, where traditional cameras struggle due to signal-to-noise ratio limitations and motion artifacts. However, exploiting event cameras for LLIE is not straightforward. Existing methods [44, 42] that rely solely on event cameras often face limitations due to the sparse nature of the event data, which can lead to poor spatial resolution and difficulty in handling noise under extreme low light conditions. Furthermore, fusion techniques combining events and images often use simplistic approaches [17, 46, 35, 30], which fail to effectively capture the complex spatial-temporal interactions between them. As a result, these methods do not fully leverage the complementary strengths of event and image data, limiting their performance in LLIE tasks.
@@
Figure 2: A challenging example from our dataset containing an extremely low-light image (a) and sparse events (b). Compared with the result from the SOTA event-guided method EvLight (c), our EvRWKV (d) not only recovers the wheel details in the dark areas (e.g., the wheel) but also preserves edge details (e.g., the white line on the floor).
We argue that a core challenge lies in designing a fusion architecture that respects the profound heterogeneity of event and image modalities while enabling meaningful synergy between them. As shown in Fig. 1 (b), prevailing methods often adopt a decoupled design [30, 20, 45, 36, 17], processing event and image streams through separate, parallel pathways. This architectural choice inherently inhibits continuous interaction, restricting feature exchange to a single terminal fusion step. Consequently, vital low-level complementary correlations are lost during independent hierarchical processing. These include transient event edges corresponding to faint image textures. By the time the features are combined, they are too abstract to benefit from this lost complementary detail, leading to underutilized complementarity, semantic misalignment, and suboptimal enhancement, as shown in Fig. 2 (c).
@@
Conversely, as shown in Fig. 1 (a), strategies [31, 35] that merge modalities at the input level struggle with a fundamental mismatch in data structure. Events are temporally dense but spatially sparse [32], whereas images offer dense spatial information, but are temporally discrete [26]. Forcing these inherently complementary representations into a common format at the outset creates a representational bottleneck, obscuring their distinct characteristics and preventing networks from effectively learning complementary relationships. This approach may also amplify noise from both sources [10, 26], degrading the fused representation before feature extraction begins. The fundamental limitation across these paradigms is the absence of a continuous adaptive interaction between the modalities throughout the feature learning process.
@@
To address this challenge, we introduce EvRWKV, a novel framework engineered for continuous cross-modal interaction through dual-domain processing. Instead of treating fusion as a single-stage process, EvRWKV establishes persistent cross-modal interactions between the RGB data and event streams at every network layer. This ensures that features from each modality dynamically inform and refine one another throughout the entire representation learning hierarchy. At the core of our approach, the Cross-RWKV module leverages the Receptance Weighted Key Value (RWKV) architecture [40] to enable fine-grained interaction between event and image features, capturing temporal dependencies and ensuring consistent spatiotemporal alignment. Complementing this, the Event Image Spectral Fusion Enhancer (EISFE) module addresses noise and structural degradation by applying an Adaptive Gaussian Filter in the frequency domain to suppress noise while preserving key structural details, and using deformable convolutions with learnable offset fields in the spatial domain to align and fuse high-frequency details from both modalities. By ensuring dynamic interaction across all stages of feature extraction—from low-level edges and textures, which are calibrated in real time, to mid-level structures that compensate for each other’s deficiencies, and high-level abstract features that are optimized for consistency—EvRWKV achieves seamless cross-modal fusion. This end-to-end collaboration maintains continuous feature alignment, preventing early information loss or late-stage disconnection, thus delivering robust LLIE that effectively preserves fine details, ensures global consistency, and suppresses noise and structural degradation.
@@
  • We propose EvRWKV, a novel framework that establishes continuous cross-modal interaction through dual-domain processing for LLIE, establishing intrinsic information exchange between event and image modalities.
  • We present Cross-RWKV, an RWKV-based backbone that captures temporal dynamics and performs cross-modal fusion through simultaneous event-image processing for enhanced low-light performance.
  • We develop EISFE, a module operating in frequency and spatial domains to achieve noise mitigation, structure enhancement, and multi-modal feature alignment.
@@ II-A Frame-based LLIE
LLIE aims to improve the visibility and quality of images captured in dimly lit environments. Traditional frame-based LLIE methods primarily operate on standard image frames. These methods can be broadly categorized into histogram equalization [2, 1, 4], Gama Correction [16, 41] and Retinex-based algorithms [12, 54, 9, 58, 33]. The first two methods directly enhance the intensity and contrast of low light image, while the Retinex-based algorithm models the image as a combination of illumination and reflectivity. Although effective in improving visibility, these approaches often encounter limitations in handling complex lighting conditions, especially in scenes with varying illumination or dark regions. Moreover, these methods may amplify noise, particularly in low-light regions, leading to undesirable artifacts.
Recent advancements have been propelled by deep learning, particularly with convolutional neural networks (CNNs) [9, 28, 5] and Transformer architectures [55, 3, 60], which have led to the development of more sophisticated frame-based LLIE methods. Chen et al. [5] established a low light image dataset and proposed a full convolution network for enhancement. Wang et al. [48] proposed to enhance underexposed photos by learning the illumination map. Wang et al. [47] collect SDSD dataset by using mechatronic system and proposed a framework integrating progressive alignment and Retinex-based illumination prediction. Cai et al. [3] propose a one-stage Transformer framework Retinexformer for LLIE, leveraging Retinex theory and illumination-guided attention to suppress noise. However, these frame-only methods fundamentally suffer from information loss in dark regions due to sensor noise floor limitations, leading to irreversible texture degradation and motion blur amplification.
@@ II-B Event-guided LLIE
The emergence of event cameras has significantly revolutionized low-light perception by offering a new paradigm in image sensing. Unlike traditional frame-based cameras that capture images at fixed intervals, event cameras detect changes in the scene asynchronously, producing a continuous stream of events that encode pixel-level intensity changes with high temporal resolution. Rebecq et al. [42] propose a recurrent network to reconstruct high speed and high dynamic range videos from event camera data. Event-based approaches for LLE like NER-Net [34] leveraged event streams’ temporal continuity to address non-uniform illumination via learnable timestamp calibration. Subsequent works explored hybrid event-frame fusion strategies. Jiang et al. [17] established joint feature learning through attention, while [30] introduced the SDE dataset with precise spatial-temporal alignment and proposed SNR-guided feature selection in EvLight framework. Kim et al. [20] propose an event-guided end-to-end framework for joint low-light video enhancement and deblurring on the real-world RELED dataset, utilizing temporal alignment and cross-modal spectral filtering. However, current fusion architectures are still limited by methods that rely on basic fusion strategies, which fail to effectively separate modality-specific noise patterns. Additionally, many of these approaches suffer from inflexible fusion operators that lack adaptive cross-modal interaction mechanisms. Furthermore, they often underutilize the event streams’ intrinsic motion information, hindering comprehensive dynamic scene modeling.
@@ II-C Receptance Weighted Key Value
The Receptance Weighted Key Value (RWKV) architecture [40] emerges as a transformative sequential modeling paradigm, synergizing the parallel computation advantages of transformers with the memory efficiency of recurrent networks. Departing from conventional attention mechanisms constrained by quadratic complexity, RWKV introduces time-shifted receptance gates and exponentially decaying key projections to maintain hidden states [39], enabling linear computational scaling with sequence length. In computer vision, RWKV has demonstrated remarkable versatility across diverse tasks [29]. For instance, Restore-RWKV [57] employs an omnidirectional token translation layer and recurrent WKV attention to restore medical images [56] by capturing spatial dependencies, while PointRWKV [13] leverages linear complexity and multi-head matrix states to enhance geometric feature extraction in 3D point cloud processing. Vision-RWKV [8] further extends its capabilities to high-resolution image analysis by reducing spatial aggregation complexity, outperforming traditional Vision Transformers (ViTs). In real-time video analytics, the TLS-RWKV framework [61] achieves efficient temporal pattern recognition for action detection. StyleRWKV [7] utilizes RWKV models for efficient style transfer, enhancing global and local context while maintaining low memory usage and linear time complexity.
Despite these advances, RWKV’s application to cross-modal fusion scenarios—particularly for processing asynchronous, heterogeneous data streams like images and events—remains underdeveloped. Existing RWKV variants struggle with asynchronous temporal alignment and spatial-temporal sparsity gaps. While recent works like RWKV-CLIP [11] explore vision-language alignment, they primarily focus on static multimodal interactions. Our work proposes a novel perspective: establishing continuous event-image synergy, where events resolve temporal ambiguities in blurry regions through high-frequency temporal updates, while images anchor sparse event data to structural priors via spatial-semantic grounding. This approach aligns with RWKV’s inherent strength in modeling long-range dependencies while mitigating its limitations in handling asynchronous modalities, opening new pathways for applications in motion deblurring, LLIE, and neuromorphic vision systems.
@@
Figure 3: An overview of our framework. Our method consists of three parts: (a) Feature Initialization(Sec III-A), (b) Cross-modal Feature Restoration(Sec III-B), and (c) Frequency-aware Feature Fusion(Sec III-C). Specifically, Cross-modal Feature Restoration contains multiple Cross-RWKV blocks for feature alignment, and Frequency-aware Feature Fusion integrates image and event features for final output.
@@ III Methodology
We propose EvRWKV, a novel framework designed for continuous cross-modal interaction through dual-domain processin that leverages the complementary advantages of RGB data and event streams to achieve robust illumination enhancement and noise suppression. As illustrated in Fig. 3, our framework comprises three key parts: 1) Feature Initialization, 2) Cross-modal Feature Restoration, 3) Frequency-aware Feature Fusion. The framework takes a low-light image $I$ and paired event stream $\{e_{k}\}_{k=1}^{N}$ as inputs, producing an enhanced image $I_{en}$. At first, a series of convolution layers extract features from low light images and event data, and then process them through the 4-level U-shaped encoder-decoder architecture composed of cross-RWKV composed of spatial and channel hybrid components. The output of the architecture passes the EISFE module is enhanced in frequency domain and spatial domain fusion processing to improve image quality.

### v3

@@ Abstract
Event cameras offer significant potential for Low-light Image Enhancement (LLIE), yet existing fusion approaches are constrained by a fundamental dilemma: early fusion struggles with modality heterogeneity, while late fusion severs crucial feature correlations. To address these limitations, we propose EvRWKV, a novel framework that enables continuous cross-modal interaction through dual-domain processing, which mainly includes a Cross-RWKV Module to capture fine-grained temporal and cross-modal dependencies, and an Event Image Spectral Fusion Enhancer (EISFE) module to perform joint adaptive frequency-domain denoising and spatial-domain alignment. This continuous interaction maintains feature consistency from low-level textures to high-level semantics. Extensive experiments on the real-world SDE and SDSD datasets demonstrate that EvRWKV significantly outperforms only image-based methods by 1.79 dB and 1.85 dB in PSNR, respectively. To further validate the practical utility of our method for downstream applications, we evaluated its impact on semantic segmentation. Experiments demonstrate that images enhanced by EvRWKV lead to a significant 35.44% improvement in mIoU.
@@ I Introduction
Capturing high-quality visual content under low-light conditions is a critical challenge in computer vision [li2021low1]. Images acquired in such environments suffer from severe noise and underexposure [li2021low1], degrading downstream applications. While traditional frame-based low-light image enhancement (LLIE) methods have made progress [lamba2021restoring, wang2021seeing, fu2016weighted, wang2023low], they are often constrained by a fundamental trade-off. Attempts to brighten the image can severely amplify noise, while models trained on synthetic data often fail to generalize to the diverse and unpredictable nature of real-world scenes [xu2022snr, cai2023retinexformer, wang2023ultra, 10810469, li2021low].
@@
Event cameras, using bio-inspired sensors [lichtsteiner2008128], are widely used in the field of computer vision due to their high dynamic range (HDR) and microsecond level time resolution [scheerlinck2019ced, zheng2023deep], such as image restoration [jiang2023event], video super-resolution [lu2023learning], optical flow [luo2023learning], deblurring [jiang2020learning, kim2022event, chen2025event], etc., and emerge as a powerful solution for LLIE. By asynchronously capturing per-pixel brightness changes, events excel at preserving structural details even in extreme darkness and are less susceptible to motion blur. However, exploiting this synergy for LLIE is not straightforward. Methods [stoffregen2020reducing, rebecq2019high] relying solely on event cameras are inherently limited by data sparsity, leading to poor spatial resolution and noise artifacts. Furthermore, fusion techniques combining events and images often use simplistic approaches [jiang2023event, wang2020event, liu2023low, liang2024towards], which fail to effectively capture the complex spatial-temporal interactions between them. As a result, these methods do not fully leverage the complementary strengths of event and image data, limiting their performance in LLIE tasks.
@@
Figure 2: A challenging example from our dataset containing an extremely low-light image (a) and sparse events (b). Compared with the result from the state-of-the-art event-guided method EvLight (c), our EvRWKV (d) not only recovers the wheel details in the dark areas (e.g., the wheel) but also preserves edge details (e.g., the white line on the floor).
@@
The core challenge lies in navigating the fundamental dilemma posed by the two dominant fusion paradigms. As shown in Fig. 1 (b), the prevailing methods often adopt a decoupled design [liang2024towards, kim2024towards, sun2025low, liu2025bidirectional, jiang2023event], processing the events and images through separate, parallel pathways. This architectural choice inherently inhibits continuous interaction, restricting feature exchange to a single terminal fusion step. Consequently, vital low-level complementary correlations are lost during independent hierarchical processing. These include transient event edges corresponding to faint image textures. By the time the features are combined, they are too abstract to benefit from this lost complementary detail, leading to underutilized complementarity and suboptimal enhancement, as shown in Fig. 2 (c). Conversely, as shown in Fig. 1 (a), strategies [liang2023coherent, liu2023low] that merge modalities at the input level struggle with a fundamental mismatch in data structure. Events are temporally dense but spatially sparse [lichtsteiner2008128], whereas images offer dense spatial information, but are temporally discrete [li2021low1]. Forcing these inherently complementary representations into a common format at the outset creates a representational bottleneck, obscuring their distinct characteristics and preventing the network from effectively learning their synergistic relationship. This approach may also amplify noise from both sources [gallego2020event, li2021low1], degrading the fused representation before feature extraction begins. Ultimately, the fundamental limitation across both paradigms is the absence of a continuous adaptive interaction between the modalities throughout the entire feature learning process.
@@
To address this challenge, we introduce EvRWKV, a novel framework engineered for continuous cross-modal interaction through a dual-domain processing strategy. Instead of treating fusion as a single, discrete event, EvRWKV establishes a persistent dialogue between the image and event streams throughout the entire network hierarchy. At the core of our approach, the Cross-RWKV Module leverages the Receptance Weighted Key Value (RWKV) architecture [peng2023rwkv] to enable fine-grained interaction between event and image features and ensuring consistent spatiotemporal alignment. Complementing this, the Event Image Spectral Fusion Enhancer (EISFE) module operates in a dual domain to jointly suppress noise, which is a critical issue in early fusion, while ensuring precise alignment of complementary structural details. By ensuring dynamic interaction across all stages of feature extraction, spanning from low-level textures to high-level semantics, EvRWKV achieves a seamless and holistic cross-modal fusion. By maintaining continuous feature alignment, this end-to-end collaboration prevents early information loss and late-stage disconnection, delivering a robust LLIE that effectively preserves fine details and suppresses noise and artifacts.
@@
  • We propose EvRWKV, a novel framework that establishes continuous cross-modal interaction through dual-domain processing for LLIE, effectively overcoming the limitations of both early and late fusion paradigms for robust low-light enhancement.
  • We design Cross-RWKV, an RWKV-based backbone that maintains feature consistency from low-level to high-level to preserve and leverage cross-modal correlations.
  • We introduce EISFE, a dual-domain module that performs joint noise mitigation and feature alignment, ensuring a robust and clean fusion process.
@@ II-A Frame-based LLIE
LLIE aims to improve the visibility and quality of images captured in dimly lit environments. Traditional frame-based LLIE methods primarily operate on standard image frames. These methods can be broadly categorized into histogram equalization [arici2009histogram, abdullah2007dynamic, celik2011contextual], Gama correction [huang2012efficient, rahman2016adaptive] and Retinex-based algorithms [guo2016lime, xu2020star, qiang2025gwretinex, zhao2021retinexdip, 10810469]. The first two methods directly enhance the intensity and contrast of low-light image, while the Retinex-based algorithm models the image as a combination of illumination and reflectivity. Although effective in improving visibility, these approaches often encounter limitations in handling complex lighting conditions, especially in scenes with varying illumination or dark regions. Moreover, these methods may amplify noise, particularly in low-light regions, leading to undesirable artifacts.
Recent advancements have been propelled by deep learning, particularly with convolutional neural networks (CNNs) [fu2016weighted, li2018structure, chen2018learning] and Transformer architectures [xu2022snr, cai2023retinexformer, zhu2024ghost], which have led to the development of more sophisticated frame-based LLIE methods. Chen et al. [chen2018learning] established a low light image dataset and proposed a full convolution network for enhancement. Wang et al. [wang2019underexposed] proposed to enhance underexposed photos by learning the illumination map. Wang et al. [wang2021seeing] collect SDSD dataset by using mechatronic system and proposed a framework integrating progressive alignment and Retinex-based illumination prediction. Cai et al. [cai2023retinexformer] propose a one-stage Transformer framework Retinexformer for LLIE, leveraging Retinex theory and illumination-guided attention to suppress noise. However, these frame-only methods fundamentally suffer from information loss in dark regions due to sensor noise floor limitations, leading to irreversible texture degradation and motion blur amplification.
@@ II-B Event-guided LLIE
The emergence of event cameras has significantly revolutionized low-light perception by offering a new paradigm in image sensing. Unlike traditional frame-based cameras that capture images at fixed intervals, event cameras detect changes in the scene asynchronously, producing a continuous stream of events that encode pixel-level intensity changes with high temporal resolution. Rebecq et al. [rebecq2019high] propose a recurrent network to reconstruct high speed and high dynamic range videos from event camera data. Event-based approaches for LLIE like NER-Net [liu2024seeing] leveraged event streams’ temporal continuity to address non-uniform illumination via learnable timestamp calibration. Subsequent works explored hybrid event-frame fusion strategies. Jiang et al. [jiang2023event] established joint feature learning through attention, while [liang2024towards] introduced the SDE dataset with precise spatial-temporal alignment and proposed SNR-guided feature selection in EvLight framework. Recent advances in sensor fusion enhancement [ragavendirane2025low] have demonstrated improved cross-modal integration through adaptive weighting mechanisms. Kim et al. [kim2024towards] propose an event-guided end-to-end framework for joint low-light video enhancement and deblurring on the real-world RELED dataset, utilizing temporal alignment and cross-modal spectral filtering. Additionally, prior-guided approaches [wang2025low] have shown promise in leveraging temporal consistency and historical information to stabilize enhancement under dynamic illumination. However, current fusion architectures are still limited by methods that rely on basic fusion strategies, which fail to effectively separate modality-specific noise patterns. Additionally, many of these approaches suffer from inflexible fusion operators that lack adaptive cross-modal interaction mechanisms. Furthermore, they often underutilize the event streams’ intrinsic motion information, hindering comprehensive dynamic scene modeling.
@@ II-C Receptance Weighted Key Value
The Receptance Weighted Key Value (RWKV) architecture [peng2023rwkv] emerges as a transformative sequential modeling paradigm, synergizing the parallel computation advantages of transformers with the memory efficiency of recurrent networks. Departing from conventional attention mechanisms constrained by quadratic complexity, RWKV introduces time-shifted receptance gates and exponentially decaying key projections to maintain hidden states [peng2024eagle], enabling linear computational scaling with sequence length. In computer vision, RWKV has demonstrated remarkable versatility across diverse tasks [li2024survey]. For instance, Restore-RWKV [yang2025restore] employs an omnidirectional token translation layer and recurrent WKV attention to restore medical images [yang2025explicit] by capturing spatial dependencies, while PointRWKV [he2025pointrwkv] leverages linear complexity and multi-head matrix states to enhance geometric feature extraction in 3D point cloud processing. Vision-RWKV [duan2024vision] further extends its capabilities to high-resolution image analysis by reducing spatial aggregation complexity, outperforming traditional Vision Transformers (ViTs). In real-time video analytics, the TLS-RWKV framework [zhu2024tls] achieves efficient temporal pattern recognition for action detection. StyleRWKV [dai2024stylerwkv] utilizes RWKV models for efficient style transfer, enhancing global and local context while maintaining low memory usage and linear time complexity.
Despite these advances, the application of RWKV [peng2023rwkv] to cross-modal fusion with asynchronous data streams remains largely unexplored. Effectively modeling such data requires addressing the dual challenges of temporal sparsity and asynchronicity [gallego2020event]. While Transformers’ all-to-all attention [vaswani2017attention] is powerful, this un-differentiated global modeling can create spurious temporal dependencies by wrongly associating distant, unrelated events. On the other hand, conventional RNNs, with their data-dependent gating mechanisms, often exhibit training instability and struggle to capture long-range dependencies [bengio1994learning] in highly sparse and noisy event data. The RWKV architecture is uniquely positioned to overcome these limitations. Its principled, non-learned time-decay mechanism provides a robust inductive bias, ensuring stable memory retention that aligns with the physical prior of a decaying event influence. Furthermore, its linear complexity and recurrent nature enable the native processing of long, continuous sequences, thereby preserving their asynchronous structure. We propose a novel cross-modal synergy wherein the stable, temporally-aware backbone of RWKV allows events to resolve image ambiguities while images ground sparse event data, opening new pathways for applications in motion deblurring and LLIE.
@@
Figure 3: Overall Architecture of the proposed EvRWKV. Our method consists of three parts: (a) Feature Initialization(Sec III-A), (b) Cross-modal Feature Restoration(Sec III-B), and (c) Frequency-aware Feature Fusion(Sec III-C). Specifically, Cross-modal Feature Restoration contains multiple Cross-RWKV blocks for feature alignment, and Frequency-aware Feature Fusion integrates image and event features for final output.
@@ III Methodology
We propose EvRWKV, a novel framework designed for continuous cross-modal interaction through dual-domain processin that leverages the complementary advantages of RGB data and event streams to achieve robust illumination enhancement and noise suppression. As shown in Fig. 3, our framework comprises three key parts: 1) Feature Initialization, 2) Cross-modal Feature Restoration, 3) Frequency-aware Feature Fusion. The framework takes a low-light image $I$ and paired event stream $\{e_{k}\}_{k=1}^{N}$ as inputs, producing an enhanced image $I_{en}$. At first, a series of convolution layers extract features from low light images and event data, and then process them through the 4-level U-shaped encoder-decoder architecture composed of cross-RWKV composed of spatial and channel hybrid components. The output of the architecture passes the EISFE module is enhanced in frequency domain and spatial domain fusion processing to improve image quality.